# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ B.5.3 Bounding the interpolation operator via Fourier transform
Furthermore, recall from the definition of (70) that $P(f)(2z)=f(z),\;\forall z\in\mathbb{Z}^{d}$. Hence, for all $z\in\mathbb{Z}^{d}$, there exists a $z^{\prime}\in\mathbb{Z}^{d}$ satisfying $\|z-z^{\prime}\|_{2}\leq\sqrt{d}$, such that $P(f)(z^{\prime})=f(z^{\prime}/2)$. Therefore, the supremum norm of $P(f)$ can be controlled as
$\|P(f)\|_{\infty}\leq\|f\|_{\infty}+2\pi\sqrt{d}\|\Xi(\mathcal{F}(P(f)))\|_{1}=\|f\|_{\infty}+2\pi\sqrt{d}\|\Xi(M(\mathcal{F}(f)))\|_{1}.$ (84)
Note that assertion (84) provides a convenient way for controlling the supremum norm of $P^{j}(f)$ with the help of the $\ell^{1}$ norm of $\Xi(M^{j}(\mathcal{F}(f)))$. However, the operator $M$ is not Markov and hence the arguments applied in Case 1 do not go through. Therefore, we proceed by defining an auxiliary operator $\tilde{M}$ such that for all $\rho\in\ell^{1}(\mathbb{T}([0,1]^{d}))$ satisfying $\rho(k)=0,k\in\mathbb{K}_{0}$,
$\tilde{M}(\rho)(\xi)=\frac{\|\xi\|_{\ell^{1}}}{\|2\xi\|_{\ell^{1}}}\rho(2\xi)|\psi(\xi)|,$ (85)
where the function is defined as its limit if $\|2\xi\|_{\ell^{1}}=0$. Note that
$\tilde{M}(\Xi(\rho))(\xi)=\Xi(M(\rho))(\xi).$ (86)
Note that in contrast to the operator $M$ representing the Fourier transform of the polynomial interpolation with the Fourier transform of the function, the operator $\tilde{M}$ acts on $\Xi(\rho)$, which is associated with the gradient of the function.
Then, by applying (84) and (86) recursively, we obtain for all $j\in\mathbb{N}$,
$\|P^{j}(f)\|_{\infty}\leq\|f\|_{\infty}+2\pi\sqrt{d}\sum_{j^{\prime}=1}^{j}\|\tilde{M}^{j^{\prime}}(\Xi(\mathcal{F}(f)))\|_{1}.$ (87)
Hence, instead of bounding the norm of $M^{j}\mathcal{F}(f)$ as in Case 1, we need to bounding the sum of the norms of $\tilde{M}^{j}\Xi(\mathcal{F}(f)))$. Then, similarly to the definition of $\mathcal{K}_{M}$ in (82), for all $A=(a_{1},a_{2})^{d}\subset\mathbb{T}([0,1]^{d})$, let
$\mathcal{K}_{\tilde{M}}(\xi_{1},A)=\frac{1}{2^{d}}\sum_{\xi_{2}\in\mathbb{T}([0,1]^{d}):2\xi_{2}=\xi_{1}}\frac{\|\xi_{2}\|_{\ell^{1}}}{\|\xi_{1}\|_{\ell^{1}}}|\psi(\xi_{2})|\mathbbm{1}_{\xi_{2}\in A}.$
Then, for any non-negative function $\rho\in\ell^{1}(\mathbb{T}([0,1]^{d}))$,
$\int_{\mathbb{T}([0,1]^{d})}\tilde{M}(\rho)(\xi)d\xi=\int_{\mathbb{T}([0,1]^{d})}\rho(\xi)\mathcal{K}_{\tilde{M}}(\xi,\mathbb{T}([0,1]^{d}))d\xi.$ (88)
Note, that for all $\xi\in\mathbb{T}([0,1]^{d})$, $\mathcal{K}_{\tilde{M}}(\xi,\cdot)$ is a nonnegative finite measure.
Next assume that the measure of the whole space is bounded away from one, i.e. $\exists c\in(0,1)$, such that
$\sup_{\xi\in\mathbb{T}([0,1]^{d})}\mathcal{K}_{\tilde{M}}(\xi,\mathbb{T}([0,1]^{d}))\leq 1-c.$ (89)
Then, in view of $\Xi(\rho)\geq 0$, for all $f\in\ell^{1}({\mathbb{Z}^{d}})$,
@@
$\leq$ $\int_{\mathbb{T}([0,1]^{d})}\Xi(\mathcal{F}(f))(\xi)(1-c)d\xi=(1-c)\|\Xi(\mathcal{F}(f))\|_{1}.$ (90)
Plugging (B.5.3) into the upper bound (87), results in
@@
proving that the right hand side of (B.5.2) is bounded by a constant and therefore verifying the statement of the lemma.
Therefore, it remains to verify the upper bound (89), where the LHS is a function on $\mathbb{T}([0,1]^{d})$. The analytical derivation of the above upper bound is highly challenging. However, note, that it is a uniformly continuous function of $\xi$, hence numerical evaluation of this function can provide the required upper bound for arbitrary combination of the smoothness $\alpha$ and the dimension $d$. The results for $d=1,2$ and various smoothness $\alpha$ are displayed in Figure 8. One can observe that for all regularity parameters in our statement, the constant $c$ in equation (89) is bound away from $1$. Our approach has the potential to be generalized to arbitrary combinations of $\alpha$ and $d$. However, the operator $\tilde{M}$ defined in (85) may no longer be sufficient to guarantee that the maximum transition measure is no greater than $1$. It is possible to extend the torus $\ell^{1}$ norm to some other functions based on $\alpha,d$ and redefine an operator using this function, but such procedure would further increase the level of technicality and hence is beyond the scope of the present paper.
Figure 8: Evaluation of LHS of equation (89) for $d=1,2$ and various regularity parameters.
B.6 Proof of Lemma 7
@@
Recall the definition of $G_{j}$ in (33), the operator $G_{j}$ is dependent on $\alpha$ and $j$. We discuss based on the values of $\alpha,j_{1}$ and $j_{2}$.
Case 1 First, let us consider the case $\tau\gamma^{-j_{2}}\geq 1$. The key idea is that as the distance between $X_{i}$ and $\mathrm{pa}(X_{i})$ is sufficiently large, the Gaussian interpolation will be sufficiently small.
@@
Since $K$ is the Matérn covariance function, we can write $K_{x_{1},x_{2}}=K_{0}(\|x_{1}-x_{2}\|_{2})$, where $K_{0}(x)$ decays to zero as $x\to\infty$. Thus, for $\tau\gamma^{-j}$ sufficiently large, $K_{\tau\gamma^{-j}A,\tau\gamma^{-j}A}$ is a diagonal dominating matrix and we can bound its minimal eigenvalue using Gershgorin circle theorem as
@@
Upper bounding each element of the vector, we have the following $\ell^{2}$ norm control for the covariance between $X_{i}$ and $\mathrm{pa}(X_{i})$:
@@
As $\tau\gamma^{-j}\to\infty$, $K_{0}(\tau\gamma^{-j}c_{d})$ and $K_{0}(\tau\gamma^{-j+1}c_{d})$ tend to zero and therefore the above display also goes to zero. Thus, there exists a constant $c^{\prime}$ independent of $j$, such that if $\tau\gamma^{-j}\geq c^{\prime}$, $1-(m-1)K_{0}(\tau\gamma^{-j}c_{d})]^{-1}\sqrt{m}K_{0}(\tau\gamma^{-j+1}c_{d})\leq 1/\sqrt{m}$ and $G_{j}(f)(X)\leq\sup_{X^{\prime}\in\mathrm{pa}(X)}|f(X^{\prime})|$. Since this holds for arbitrary $X_{i}\in\mathcal{N}_{j+1}$, in view of the definition of $G_{j}$ in (33), this proves $\|G_{j}\|_{\infty}\leq 1$. If $\tau\gamma^{-j_{2}}\geq c^{\prime}$, this proves the lemma. Otherwise, if $1\leq\tau\gamma^{-j_{2}}<c^{\prime}$, let us consider the index set
@@
Then the cardinality of $\mathcal{J}$ is bounded by an absolute constant $\log_{\gamma}c^{\prime}$. For all $j\in\mathcal{J}$, we use the naive norm bound that
@@
resulting in $\|G_{j}\|_{\infty}\leq m$. For all $j_{1}\leq j\leq j_{2}$ and $j\not\in\mathcal{J}$, we still have $\|G_{j}\|_{\infty}\leq 1$, thus
$\|G_{j_{2}}\cdots G_{j_{1}+1}G_{j_{1}}\|_{\infty}\leq m^{\log_{\gamma}c^{\prime}}.$
Case 2 Next, we assume that $\tau\gamma^{-j_{1}}\leq 1$, in which case $\tau\gamma^{-j}<1,\;\forall j_{1}\leq j\leq j_{2}$. The idea is that as elements in $\mathrm{pa}(X_{i})\cup\{X_{i}\}$ get very close to each other, the flat limit phenomenon of Lemma 6 occurs and Gaussian interpolation can well approximated by polynomial interpolations.
For all $j\in\mathbb{N}$, let us denote by $E_{j}=G_{j}-P_{j}$ the difference between Gaussian and polynomial interpolation at layer $j$. By Condition 2, for all $X_{i}\in\mathcal{N}_{j}$, $X_{i}$ and its parent set is included in a cube $\mathcal{C}$ with side length no greater than $c_{L}\gamma^{-j}$, which can be regarded as rescaling the unit cube $0,1]^{d}$ with the rescaling parameter $c_{L}\gamma^{-j}$. Combining this with Lemma 6, we have
$\|E_{j}\|_{\infty}\lesssim\gamma^{-2j(\alpha-\underline{\alpha})}+\gamma^{-j}.$
The key idea is to show that the operator $E_{j}$ tends to zero exponentially fast, which implies that $\sum_{j=1}^{\infty}\|E_{j}\|_{\infty}$ is finite. Then we express the recursive Gaussian interpolation operator with the help of the polynomial interpolation operator and the error term. The previously discussed summability will allow us to control the terms in the upper bound.
Let $\alpha^{*}=\min\{2(\alpha-\underline{\alpha}),1\}$ and note that $\|E_{j}\|_{\infty}\lesssim\gamma^{-\alpha^{*}j}$. For $j=j_{1}$, write
$G_{j_{1}}=P_{j_{1}}+E_{j_{1}}=a_{j_{1}}P_{j_{1}}H_{j_{1}}+a_{j_{1}+1}H_{j_{1}+1}$ (91)
@@
$\|H_{j_{1}+1}\|_{\infty}=\|E_{j_{1}}/\gamma^{-j_{1}\alpha^{*}}\|_{\infty}\lesssim 1.$
We prove below that for all $j_{1}\leq j\leq j_{2}+1$, there exists an operator $\|H_{j+1}\|_{\infty}\lesssim 1$ and a constant $a_{j}\in[0,+\infty)$, such that
$G_{j}\cdots G_{j_{1}+1}G_{j_{1}}=\sum_{j^{\prime}=j_{1}}^{j}a_{j^{\prime}}P_{j}\cdots P_{j^{\prime}+1}P_{j^{\prime}}H_{j^{\prime}}+a_{j+1}H_{j+1}.$ (92)
@@
Then, let us introduce the notation $S_{j}=\sum_{j^{\prime}=j_{1}}^{j}a_{j^{\prime}}$, and note that by the preceding display
@@
for some universal constant $c^{\prime}>0$. Therefore, recalling the definition of $\tilde{\vartheta}_{n}$
$\|G_{j_{2}}\cdots G_{j_{1}+1}G_{j_{1}}\|_{\infty}\leq\vartheta_{n}S_{j_{2}}\leq\vartheta_{n}\exp\left[\frac{c^{\prime}\vartheta_{n}}{1-\gamma^{-\alpha^{*}}}\right]=\tilde{\vartheta}_{n},$
@@
Hence it remains to prove (92). We proceed with proof by induction. By equation (91), the claim holds for $j=j_{1}$. Now suppose the claim holds for $j=l$. Then for $j=l+1$, we have
@@
then obviously $\|H_{l+2}\|\leq 1$ Recalling the definition of $\vartheta_{n}$ and the induction assumption that $\|H_{j}\|_{1}\leq 1,\forall j_{1}\leq j\leq l+1$, we have
@@
Case 3 Finally, if $\tau\gamma^{-j_{2}}<1<\tau\gamma^{-j_{1}}$, then there exists $j_{1}<j_{3}\leq j_{2}$, such that $\tau\gamma^{-j_{2}}\leq\tau\gamma^{-j_{3}}\leq 1<\tau\gamma^{-j_{3}+1}\leq\gamma^{-j_{1}}$. For $j_{1}\leq j\leq j_{3}-1$, we can apply Case 2, whereas for $j_{3}\leq j\leq j_{2}$, we can apply Case 1. Combining both parts concludes this case. ∎
C Proofs of the lemmas and theorems in Section 5.3 and 5.4
C.1 Proof of Theorem 2

### v4

@@ E.5.3 Bounding the interpolation operator via Fourier transform
Furthermore, recall from the definition of (E.24) that $P(f)(2z)=f(z),\;\forall z\in\mathbb{Z}^{d}$. Hence, for all $z\in\mathbb{Z}^{d}$, there exists a $z^{\prime}\in\mathbb{Z}^{d}$ satisfying $\|z-z^{\prime}\|_{2}\leq\sqrt{d}$, such that $P(f)(z^{\prime})=f(z^{\prime}/2)$. Therefore, the supremum norm of $P(f)$ can be controlled as
$\|P(f)\|_{\infty}\leq\|f\|_{\infty}+2\pi\sqrt{d}\|\Xi(\mathcal{F}(P(f)))\|_{1}=\|f\|_{\infty}+2\pi\sqrt{d}\|\Xi(M(\mathcal{F}(f)))\|_{1}.$ (E.38)
Note that assertion (E.38) provides a convenient way for controlling the supremum norm of $P^{j}(f)$ with the help of the $L_{1}$ norm of $\Xi(M^{j}(\mathcal{F}(f)))$. However, the operator $M$ is not Markov and hence the arguments applied in Case 1 do not go through. Therefore, we define an auxiliary operator $\tilde{M}$, such that, for all $\rho\in L_{1}(\mathbb{T}([0,1]^{d}))$ satisfying $\rho(k)=0,k\in\mathbb{K}_{0}$,
$\tilde{M}(\rho)(\xi)=\frac{\|\xi\|_{L_{1}}}{\|2\xi\|_{L_{1}}}\rho(2\xi)|\psi(\xi)|,$ (E.39)
where the function is interpreted as its limit when $\|2\xi\|_{L_{1}}=0$. It is easy to see that
$\tilde{M}(\Xi(\rho))(\xi)=\Xi(M(\rho))(\xi).$ (E.40)
Note that in contrast to the operator $M$, which maps the Fourier transform of the function to the Fourier transform of its polynomial interpolation, the operator $\tilde{M}$ acts on $\Xi(\rho)$, which is associated with the gradient of the function.
Then, by applying (E.38) and (E.40) recursively, we obtain for all $j\in\mathbb{N}$,
$\|P^{j}(f)\|_{\infty}\leq\|f\|_{\infty}+2\pi\sqrt{d}\sum_{j^{\prime}=1}^{j}\|\tilde{M}^{j^{\prime}}(\Xi(\mathcal{F}(f)))\|_{1}.$ (E.41)
Hence, instead of bounding the norm of $M^{j}\mathcal{F}(f)$ as in Case 1, we need to bound the sum of the norms of $\tilde{M}^{j}\Xi(\mathcal{F}(f))$. Then, similarly to the definition of $\mathcal{K}_{M}$ in (E.36), for all $A=(a_{1},a_{2})^{d}\subset\mathbb{T}([0,1]^{d})$, let
$\mathcal{K}_{\tilde{M}}(\xi_{1},A)=\frac{1}{2^{d}}\sum_{\xi_{2}\in\mathbb{T}([0,1]^{d}):2\xi_{2}=\xi_{1}}\frac{\|\xi_{2}\|_{L_{1}}}{\|\xi_{1}\|_{L_{1}}}|\psi(\xi_{2})|\mathbbm{1}_{\xi_{2}\in A}.$
Then, for any non-negative function $\rho\in L_{1}(\mathbb{T}([0,1]^{d}))$,
$\int_{\mathbb{T}([0,1]^{d})}\tilde{M}(\rho)(\xi)d\xi=\int_{\mathbb{T}([0,1]^{d})}\rho(\xi)\mathcal{K}_{\tilde{M}}(\xi,\mathbb{T}([0,1]^{d}))d\xi.$ (E.42)
Note that for all $\xi\in\mathbb{T}([0,1]^{d})$, $\mathcal{K}_{\tilde{M}}(\xi,\cdot)$ is a nonnegative finite measure.
Next, assume that the measure of the whole space is bounded away from one, i.e. $\exists c\in(0,1)$, such that
$\sup_{\xi\in\mathbb{T}([0,1]^{d})}\mathcal{K}_{\tilde{M}}(\xi,\mathbb{T}([0,1]^{d}))\leq 1-c.$ (E.43)
Then, in view of $\Xi(\rho)\geq 0$, for all $f\in L_{1}({\mathbb{Z}^{d}})$,
@@
$\leq$ $\int_{\mathbb{T}([0,1]^{d})}\Xi(\mathcal{F}(f))(\xi)(1-c)d\xi=(1-c)\|\Xi(\mathcal{F}(f))\|_{1}.$ (E.44)
Plugging (E.5.3) into the upper bound (E.41), results in
@@
proving that the right-hand side of (E.5.2) is bounded by a constant and therefore verifying the statement of the lemma.
Therefore, it remains to verify the upper bound (E.43), where the LHS is a function on $\mathbb{T}([0,1]^{d})$. While the analytical derivation of this bound is highly challenging, the function is uniformly continuous in $\xi$, which allows for numerical evaluation to obtain the required upper bound for arbitrary combinations of the smoothness $\alpha$ and dimension $d$. The results for $d=1,2$ and various values of $\alpha$ are displayed in Figure E.1. One can observe that, for all regularity parameters considered, the constant $c$ in equation (E.43) is bound away from $1$. Our approach could potentially be extended to arbitrary combinations of $\alpha$ and $d$. However, in such general settings, the operator $\tilde{M}$ defined in (E.39) may no longer guarantee that the maximum transition measure does not exceed $1$. One could consider extending the torus $L_{1}$ norm to other functions depending $\alpha,d$ and redefine the operator accordingly. Nonetheless, this would introduce additional technical complexity and is beyond the scope of the present paper.
Figure E.1: Evaluation of LHS of equation (E.43) for $d=1,2$ and various regularity parameters.
E.6 Proof of Lemma 4
@@
We distinguish three cases according to the values of $j_{1}$ and $j_{2}$.
Case 1 First, consider the case $\tau\gamma^{-j_{2}}\geq 1$. The key observation is that when the distance between $X_{i}$ and $\mathrm{pa}(X_{i})$ is sufficiently large, the corresponding Gaussian interpolation term becomes negligibly small.
@@
Since $K$ is the Matérn covariance kernel, we can write $K_{x_{1},x_{2}}=K_{0}(\|x_{1}-x_{2}\|_{2})$, where $K_{0}(x)$ decays to zero as $x\to\infty$. Thus, for $\tau\gamma^{-j}$ sufficiently large, $K_{\tau\gamma^{-j}A,\tau\gamma^{-j}A}$ is diagonally dominant, and we can bound its minimal eigenvalue using Gershgorin circle theorem (Theorem 6.1.1 of [24]) as
@@
Upper bounding each element of the vector, we obtain the following $L_{2}$-norm bound for the covariance between $X_{i}$ and $\mathrm{pa}(X_{i})$:
@@
As $\tau\gamma^{-j}\to\infty$, $K_{0}(\tau\gamma^{-j}c_{d})$ and $K_{0}(\tau\gamma^{-j+1}c_{d})$ tend to zero and therefore, the above display also goes to zero. Thus, there exists a constant $c^{\prime}$ independent of $j$, such that if $\tau\gamma^{-j}\geq c^{\prime}$, $1-(m-1)K_{0}(\tau\gamma^{-j}c_{d})]^{-1}\sqrt{m}K_{0}(\tau\gamma^{-j+1}c_{d})\leq 1/\sqrt{m}$ and $G_{j}(f)(X)\leq\sup_{X^{\prime}\in\mathrm{pa}(X)}|f(X^{\prime})|$. Since this holds for arbitrary $X_{i}\in\mathcal{N}_{j+1}$, in view of the definition of $G_{j}$ in (28), $\|G_{j}\|\leq 1$. If $\tau\gamma^{-j_{2}}\geq c^{\prime}$, this proves the lemma. Otherwise, if $1\leq\tau\gamma^{-j_{2}}<c^{\prime}$, let us consider the index set
@@
Then the cardinality of $\mathcal{J}$ is bounded by an absolute constant $\log_{\gamma}c^{\prime}$. For all $j\in\mathcal{J}$, we use the norm bound
@@
resulting in $\|G_{j}\|\leq m$. For all $j_{1}\leq j\leq j_{2}$ and $j\not\in\mathcal{J}$, we still have $\|G_{j}\|\leq 1$, thus
$\|G_{j_{2}}\cdots G_{j_{1}+1}G_{j_{1}}\|\leq m^{\log_{\gamma}c^{\prime}}.$
Case 2 Next, we assume that $\tau\gamma^{-j_{1}}\leq 1$, in which case $\tau\gamma^{-j}<1,\;\forall j_{1}\leq j\leq j_{2}$. The key idea is that as elements in $\mathrm{pa}(X_{i})\cup\{X_{i}\}$ become increasingly close to each other, the flat-limit phenomenon of Lemma 3 takes effect, and the Gaussian interpolation can be well approximated by polynomial interpolations.
For all $j\in\mathbb{N}$, let us denote by $E_{j}=G_{j}-P_{j}$ the difference between Gaussian and polynomial interpolation operators at layer $j$, with its supreme operator normed defined in the same way as $P_{j}$ in Section 4.2.2. By Condition 2, for every $X_{i}\in\mathcal{N}_{j}$, the point $X_{i}$ together with its parent set is included in a cube $\mathcal{C}$, whose side length is at most $c_{L}\gamma^{-j}$. This configuration can be viewed as a rescaled version of the unit cube $0,1]^{d}$ with scaling parameter $c_{L}\gamma^{-j}$. Combining this observation with Lemma 3, we obtain
$\|E_{j}\|\lesssim\gamma^{-2j(\alpha-\underline{\alpha})}+\gamma^{-j}.$
The key idea is to show that the operator $E_{j}$ decays exponentially fast to zero as $j\rightarrow\infty$, which in particular implies that $\sum_{j=1}^{\infty}\|E_{j}\|<\infty$. With this summability at hand, we express the recursive Gaussian interpolation operator in terms of the polynomial interpolation operator together with the accumulated error contribution $E_{j}$. The finiteness of the series $\sum_{j=1}^{\infty}\|E_{j}\|$ will then allow us to control all error terms in the upper bound.
Let $\alpha^{*}=\min\{2(\alpha-\underline{\alpha}),1\}$ and note that $\|E_{j}\|\lesssim\gamma^{-\alpha^{*}j}$. For $j=j_{1}$, write
$G_{j_{1}}=P_{j_{1}}+E_{j_{1}}=a_{j_{1}}P_{j_{1}}H_{j_{1}}+a_{j_{1}+1}H_{j_{1}+1}$ (E.45)
@@
$\|H_{j_{1}+1}\|=\|E_{j_{1}}/\gamma^{-j_{1}\alpha^{*}}\|\lesssim 1.$
We prove below that for all $j_{1}\leq j\leq j_{2}+1$, there exists an operator $\|H_{j+1}\|\lesssim 1$ and a constant $a_{j}\in[0,+\infty)$, such that
$G_{j}\cdots G_{j_{1}+1}G_{j_{1}}=\sum_{j^{\prime}=j_{1}}^{j}a_{j^{\prime}}P_{j}\cdots P_{j^{\prime}+1}P_{j^{\prime}}H_{j^{\prime}}+a_{j+1}H_{j+1}.$ (E.46)
@@
Then, let us introduce the notation $S_{j}=\sum_{j^{\prime}=j_{1}}^{j}a_{j^{\prime}}$, and note that by the preceding display,
@@
for some universal constant $c^{\prime}>0$. Therefore, in view of the definition of $\tilde{\vartheta}_{n}$,
$\|G_{j_{2}}\cdots G_{j_{1}+1}G_{j_{1}}\|\leq\vartheta_{n}S_{j_{2}}\leq\vartheta_{n}\exp\left[\frac{c^{\prime}\vartheta_{n}}{1-\gamma^{-\alpha^{*}}}\right]=\tilde{\vartheta}_{n},$
@@
Hence it remains to prove (E.46). We proceed by induction. By equation (E.45), the claim holds for $j=j_{1}$. Now suppose the claim holds for $j=l$. Then, for $j=l+1$, we have
@@
then obviously $\|H_{l+2}\|\leq 1$. Recalling the definition of $\vartheta_{n}$ and the induction assumption that $\|H_{j}\|_{1}\leq 1,\forall j_{1}\leq j\leq l+1$, we have
@@
Case 3 Finally, if $\tau\gamma^{-j_{2}}<1<\tau\gamma^{-j_{1}}$, then there exists $j_{1}<j_{3}\leq j_{2}$, such that $\tau\gamma^{-j_{2}}\leq\tau\gamma^{-j_{3}}\leq 1<\tau\gamma^{-j_{3}+1}\leq\gamma^{-j_{1}}$. For $j_{1}\leq j\leq j_{3}-1$, we can apply Case 2, whereas for $j_{3}\leq j\leq j_{2}$, we can apply Case 1. Combining both parts concludes the proof. ∎
F Proofs of the lemmas and theorems in Section 4.3 and 4.4
F.1 Proof of Theorem 2